OpenStudy (anonymous):

Match these please? 3. 3x+6+1=8 4. 3x+7=8 5. 3x=1 6. x=1/3 A. Subtraction Property of Equality B. Combine like terms C. Distributive Property D. Division Property of Equality

OpenStudy (mathstudent55):

What did you do from step 1 to 2? 6 + 1 = 7. Which choice represents that?

OpenStudy (anonymous):

B?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

That is the only one i understand how to do

OpenStudy (mathstudent55):

From step 4 to 5, you went from 3x + 7 = 8 to 3x = 1. What happenned to 3x + 7 to become just 3x? What happened to the 7? What happened to 8 to become 1?

OpenStudy (anonymous):

Subtract 7 from both sides?

OpenStudy (mathstudent55):

Correct. So now match that to a choice. One is very obvious.

OpenStudy (anonymous):

4-b 5-a 6-d 3-c(not sure of this one)

OpenStudy (anonymous):

A

OpenStudy (mathstudent55):

@tiffers24 Ignore the above response, and keep working with me one step at a time.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Correct. The last one:from step 5 to step 6, you went from 3x = 1 to x = 1/3. What happened to 3x to become x and what happened to 1 to become 1/3?

OpenStudy (anonymous):

Mmm I'm not sure, divide it?

OpenStudy (anonymous):

MAth teachers here brb

OpenStudy (mathstudent55):

Right. You divide 3x by 3 to get x, and you divide 1 by 3 to get 1/3. So the last step was the division property of equality. That is choice D. The answers are: 4 B, 5 A, 6 D We can't tell what the answer to 3 is without seeing step 2, to see what happened between step 2 and 3.

OpenStudy (anonymous):

3(x+2)+1=8

OpenStudy (anonymous):

That was the first one given

OpenStudy (mathstudent55):

Then from that step to 3x + 6 + 1 = 8, the step that was done was going from 3(x + 2) to 3x + 6. That is the distributive property.

OpenStudy (anonymous):

Thank you for the help :)

OpenStudy (anonymous):

Would you happen to know the answer to this one? According to the Transitive Property of Equality, if TX = XY and XY = YZ, then TX = ___. A. TX B. XY C. YZ D. TZ

OpenStudy (mathstudent55):

The transitive property of equality is if a = b and b = c, then a = c. Look at each equality statement. At first, you don't know anything about a and c. You only know a = b amd b = c, but you can conclude a = c. Now look at your question and try to match it to a, b, and c, and see what you can conclude.

OpenStudy (anonymous):

a=b is in this question tx=xy. b=c is xy=yz so tx=yz?

OpenStudy (mathstudent55):

YZ is correct.

OpenStudy (anonymous):

Thank you very much